OpenStudy (stormswan):

PLEASE help me. I will give a medal and fan you. Which of the following is equal to square root of the cube root of 6? 6 1/6 6 1/3 6 2/3 6 3/2

OpenStudy (imstuck):

Do you have any ideas how to set this up? It helps it you know exponential forms of radicals.

OpenStudy (imstuck):

\[\sqrt{\sqrt[3]{6}}\]This is what it looks like

OpenStudy (stormswan):

yes. sorry, i don't really know how to set it up like that.

OpenStudy (imstuck):

The cube root of 6 in exponential form is\[6^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

\[\sqrt[3]{6}------ 6^{\frac{1}{3}}\]

OpenStudy (anonymous):

so \[\sqrt{6^{\frac{1}{3}}}\] can be written as \[6^{\frac{1}{3} \times \frac{1}{2}}\]

OpenStudy (anonymous):

what do you get now

OpenStudy (imstuck):

THEN you have to take that to the 1/2 power because the square root is to the 1/2 power. So this is what you have to deal with\[(6^{\frac{ 1 }{ 3 }})^{\frac{ 1 }{ 2 }}\]Do you know how to multiply exponents?

OpenStudy (stormswan):

and could you also help me on this question? Simplify square root of 5 multiplied by the cube root of 5

OpenStudy (anonymous):

why dont you try that @Stormswan

OpenStudy (stormswan):

yes, I sort of know how to.

OpenStudy (imstuck):

The rule with this is that you multiply the exponents together. 1/2 * 1/3 = 1/6 or \[6^{\frac{ 1 }{ 6 }}\]

OpenStudy (stormswan):

@thushananth01 i don't know how to simplify square and cubed root :((

OpenStudy (stormswan):

could you help me?

OpenStudy (imstuck):

square root of 5 is the same thing as\[5^{\frac{ 1 }{ 2}}\]and the cube root of 5 is the same thing as\[5^{\frac{ 1 }{ 3 }}\]When you multiply these together you follow the same rules as you would if you were multiplying \[x ^{2} \times x ^{3}\]Do you know?

OpenStudy (imstuck):

What do you do with the exponents?

OpenStudy (stormswan):

5 1/6?

OpenStudy (imstuck):

You add the exponents when you multiply their bases. You have to find a common denominator between the 2 and the 3. That would be 6. So you have, when you do that,\[5^{\frac{ 3 }{ 6 }+\frac{ 2 }{ 6 }}\]What do you get when you add those exponents?

OpenStudy (stormswan):

5 5/6 right?

OpenStudy (imstuck):

Yep! You're right! That's 5^5/6 power. Not 5 and 5/6 like a mixed fraction. That 5/6 is an exponent. Yes? You know that, right?

OpenStudy (stormswan):

yes. i know. thank you SO much! @IMStuck you are awesome!

OpenStudy (imstuck):

Thank you for that; any time you need help we are all here to help you (we do math for fun in our spare time!)

OpenStudy (stormswan):

alright! ill be sure to tag one of you next time! :)

OpenStudy (imstuck):

We'll be here!

OpenStudy (stormswan):

ok, thank you once again!